Suppose it is known that 19% of students work full-time. Suppose it is also known that 60% of the population work full-time and that 17% of the population are students. If a randomly chosen member of the population is a full-time worker, what is the probability that he/she is a student?
Answer #1

P(work full time | student) = 0.19

P(work full time) = 0.6

P(student) = 0.17

P(work full time and student) = P(work full time | student) * P(student) = 0.19 * 0.17 = 0.0323

P(student | work full time) = P(work full time and student) / P(work full time) = 0.0323 / 0.6 = 0.0538